# GastCoCo: Graph Storage and Coroutine-Based Prefetch Co-Design for Dynamic Graph Processing

### v3

@@ 3. Overview of GastCoCo
Building on the above discussion, we propose GastCoCo, an in-memory system for dynamic graph processing that employs an interleaved execution mode combining coroutines and software prefetching. GastCoCo focuses on graph storage and coroutine-based prefetch co-design, proposing a novel dynamic graph data structure friendly to prefetching techniques. Besides, GastCoCo comprises various task allocating, prefetching, and scheduling strategies to mitigate the overhead of cache misses, thereby consistently delivering robust performance for different graph processing tasks across various hardware environments.
As shown in Figure 3, GastCoCo is primarily composed of three layers: storage layer, execution layer, and adaptation layer. Given a graph processing task (i.e., graph computation or graph update), user execution begins by calling GastCoCo API, as shown in Table 1. The detailed introduction to the API is as follows:
LoadGraph(string file_path). This API is used to load graph data from a file into CBList, commonly for quickly initializing the data structure. It requires the parameter file_path to locate the file.
UpdateVertex(vertex obj_v, mode update_mode) & Update
Edge(edge obj_e, mode update_mode). These two APIs are used to update vertices and edges, respectively, and they both perform individual update operations. These updates can include insertions, deletions, or modifications, with the specific type indicated by parameter update_mode. Parameters obj_v and obj_e represent the vertex or the edge to be updated.
BatchUpdate(container update_content). This API is used to update graph data in batches. The parameter update_content is a container that holds the batch of updates, where each item includes the corresponding update data and type.
ProcessVertex(function f, bitmap active). This API is used to perform graph computation tasks that involve only vertices, such as finding the vertex with the highest degree. The parameter active is used to filter the vertices that participate in the computation and can implement the scan_vertices() (unconditional traversal), scan_vertices(cond) (conditional traversal), and read_vertex(v) (direct vertex ID comparison) definitions from Section 2.1. The parameter f is used for the user to implement specific operations on the vertices that meet the conditions.
ProcessEdge(function dense_f, function sparse_f, bit
map active). This API is used to perform graph computation tasks that involve edges. Since the computation model of GastCoCo is vertex-centric, processing edges also require active for filtering. Additionally, GastCoCo implements a hybrid push and pull computation model, so parameters dense_f and sparse_f are provided for the user to implement the corresponding edge processing functions.
GastCoCo API activates the dynamic graph processing engine, at which point the system configuration probe in the adaptation layer detects the current hardware environment to configure the system parameters of GastCoCo, e.g., numbers of coroutines per thread and hybrid prefetching strategy. To process tasks by using software prefetching via coroutine, graph data should be partitioned into several parts and allocated to each coroutine within a coroutine pool. All coroutines collaborate according to the scheduling strategy to produce the final execution results. The execution strategy tuner tailors the partition and scheduling strategy to suit the specific task to achieve better performance. The task allocator partitions the graph data according to the partition strategy and constructs the coroutine pool, while the coroutine scheduler schedules coroutines according to the scheduling strategy. Finally, the executor performs all read and write operations on the storage layer.
@@
LoadGraph (file_path) Loading graph data files.
UpdateVertex (obj_v, update_mode) Adding, deleting, and modifying vertices.
UpdateEdge (obj_e, update_mode) Adding, deleting, and modifying edges.
BatchUpdate (update_content) Handling batch update tasks.
ProcessVertex (f, active) Processing vertices task.
ProcessEdge (dense_f, sparse_f, active) Processing edges task.
Execution Layer (§5). The execution layer includes the task allocation which breaks down a graph computation task into multiple subtasks and allocates them to coroutines, the coroutine scheduler which controls the interleaving execution and prefetching, and the executor which executes the computation tasks according to the established scheduling strategy.
Adaptation Layer (§6). The adaptation layer includes two components: the system configuration probe and the execution strategy tuner. They utilize prefabricated programs to probe the optimal range of coroutine parameters (e.g., number of coroutines per thread) and the most suitable execution and prefetching strategy for the current tasks and runtime environment.
@@ 4.1. Overview
For tasks in dynamic graph processing scenarios, we have the following expectations for an ideal data structure. It should: (1) maintain cache locality as much as possible to produce efficient graph computation; (2) be easy to dynamically adjust for frequent graph updates. To enhance the performance from both perspectives simultaneously, the basic idea of our data structure design is to tailor as closely as possible to fit the hardware prefetching mechanisms and prepare for the implementation of software prefetching via coroutines.
Figure 4 presents the proposed core data structure CBList and we elaborate on its details and insights in the following. For storing vertices, we employ an ID-map table and a vertex table (array of vertex structures), as shown in Figure 4 1 and 2 to map the vertex’s original ID (any data type) to the logical ID (a unique natural number identifier that is used in graph computation). A record contains multiple fields as presented in Figure 4 3, which includes two pointers to the neighborhood (i.e., traversal pointer and update/query pointer), size (number of adjacent edges), delete flag, and the level. Here, we use a level variable to distinguish the structure of the vertex’s neighborhood. For the variable, a value 00 represents using small chunks for edge storage, while other numbers indicate using B+ trees and the value represents the number of leaf nodes in the tree.
@@ 4.2. Update-Read Balanced Edge Storage
We enumerate the advantages of our Edge Storage design as follows. Firstly, we use independent edge structures (refer to ADJ-like structures in Figure 5) for each vertex to store its neighborhood to promote parallel updates of edges and facilitate different implementations like chunks and B+ trees for different nodes. In contrast, in the CSR-like structures (e.g., PCSR (Wheatman and Xu, 2018) and Teseo (Leo and Boncz, 2021)), nodes are stored in identical and shared edge structure (refer to the CSR-like structures in Figure 5), and to update neighbors of one vertex may affect the vertex records and memory locations of other edges. Secondly, scan_edges($v_{src}$) accessed from Traversal pointer in a B+ tree can be considered as sequential data access on a linked list, which can effectively utilize hardware prefetching as discussed in subsection 2.2. Thirdly, read_edge($v_{src}$,$v_{dst}$) accessed from update/query pointer completes in $O(log(n))$.
Figure 5. CSR-like physically contiguous store (a) vs. ADJ-like mutually independent store (c) vs. prefetch-friendly logically contiguous store by adding pointers (b). E.S. means Edge Storage.
@@ 4.3. Prefecth-Friendly Global Traversal Chain
In graph computation, many important computational tasks (e.g., PageRank (Page et al., 1998)) require that all vertices and edges in the entire graph be involved in the computation. We can represent their data access patterns as a combination of scan_vertices() and scan_edges($v_{src}$). In fact, this combination equates to sequential data access over the entire graph.
@@
We have also considered the overhead of forming GTChain. One of the reasons we chose B+ trees in our Edge Storage design is the ease of forming GTChain. Each B+ tree inherently contains a traversal chain formed by leaf nodes. Its implementation for GTChain simply involves using pointers to link the leaf node chain into the global chain. However, most tree structures (i.e., balanced binary trees, B-trees, and red-black trees) require manual implementation of a traversal chain and the additional traversal chain is challenging to form as a singly linked list (each node has only one outgoing pointer) to support the jump-pointer mechanism of hardware prefetching.
@@ 5.1. Coroutine with Software Prefetching
Sequential execution and interleaved execution. This paper focuses on reducing the cache miss overhead based on the interleaving execution mode (Psaropoulos et al., 2017). Figure 6 illustrates the difference between sequential execution and interleaving execution of traversal on a toy graph in Figure 6 1. To improve the understanding of the execution process, we first demonstrate an example of scan_edges($v_{src}$) on AL (adjacency list) in Figure 6 2 for clarity: In the sequential execution mode, when the CPU accesses Node 1, the first neighbor of vertex A, a cache miss is likely to occur in the non-contiguous memory linked by pointers, making the CPU stalled while waiting for the data. On the other hand, in the interleaved execution mode, the CPU will issue a prefetch for Node 1 and switch to another task (i.e., accessing Node 3, the first neighbor of vertex B). As a result, for sequential execution, whenever the cache misses occur for accessing non-contiguous memory and pointer chasing, the CPU will experience data stalls (He et al., 2020) and must wait until the data is fetched in the cache. While for interleaving execution, each coroutine handles a linked list in charge of the computations of the corresponding task. The algorithm starts from an arbitrary task on computation. It switches to another task whenever it encounters a software prefetching instruction to fetch data (the instruction is set before accessing the next node in the linked list). When switching back to the previous task, the CPU does not need to wait for the needed data due to having loaded into the cache. Thus, the process of fetching data and computations can be overlapped and some of the time derived from cache misses due to pointer chasing can be avoided. Note that switching tasks also causes some time costs (Psaropoulos et al., 2017). So we choose C++20 stackless coroutines for interleaving and design scheduling strategy, such that the switching time is less than the time of occurring data stalls to achieve performance gains.
@@
Coroutine in Graph Computation. Next, we present the coroutine pool constructor and scheduler in graph computation. As we previously mentioned, C++20 stackless coroutines are a special kind of function and it requires manual intervention to schedule these functions. ①In algorithm 1 (algorithm 1 - algorithm 1), we present the basic logic of the coroutine pool constructor. The time complexity of the constructor is $O(m)$, where $m$ is the number of coroutines in the coroutine pool. The coroutine pool relies on the for loop (algorithm 1) to initialize each coroutine as Func. Func defines the graph access operations (as proposed in subsection 2.2) which will be detailed in algorithm 2. ②In algorithm 1 (algorithm 1 - algorithm 1), we present a pooling scheduler implementation. The overhead generated by the scheduler itself is also a cost of using coroutines. Therefore, we aim to design it as simply as possible to reduce the cost of using coroutines. Its core logic includes checking whether the coroutine pool has finished through a while loop (algorithm 1) and polling each coroutine via a for loop (algorithm 1). If a coroutine is not completed, the polling scheduler resumes execution from the suspension point (algorithm 1). If a coroutine is completed, the polling scheduler will destroy it (algorithm 1). Assuming the number of suspensions for the $i$-th coroutine in the pooling scheduler is $k_{i}$, the time complexity for the entire pooling scheduler is $O(\sum_{i=1}^{m}k_{i})$. If the tasks are evenly distributed among the coroutines, with each coroutine suspended approximately $k$ times, the overall time complexity is $O(mk)$. ③We detail the logic of edge access operations in algorithm 2. GetNeighbors($vertex$) and GetNeighbors($chain$) correspond to the scan_edges($v_{src}$) operation got edge accessing, where the difference between GetNeighbors($vertex$) and GetNeighbors
($chain$) and their usage will be elaborated in subsection 6.1. The time complexities of the two operations are $O(D)$ and $O(E_{chain})$ respectively, where $D$ is the degree of the vertex, and $E_{chain}$ is the number of edges in the chain. FindNeighbor($edge$) corresponds to read_edge($v_{src}$,$v_{dst}$). Based on CBList, regardless of whether using a binary search on small chunks or searching within a B+ tree, the time complexity for FindNeighbor($edge$) is $O(\log D)$. The user can specify the search algorithm either on small chunks or B+ tree leaf nodes (refer to algorithm 2, algorithm 2). Here, all edge access operations are constructed using coroutines. prefetch($ptr$) denotes prefetching the content pointed to by the pointer in algorithm 2, algorithm 2, and algorithm 2 (Software prefetching). In algorithm 2, algorithm 2, and algorithm 2, co_await is a keyword in C++20 and its invocation of suspend_always means that this coroutine actively entering an always suspended state. Coroutines in suspended state will be resumed by the scheduler(algorithm 1 in algorithm 1).
@@ 8 Function Scheduler($coroutine\_pool$):
15                         $coroutine\_pool[i].$destory()
@@ 5.2. Load Balancing of Coroutines
As we have described in the example of interleaving execution mode in Figure 6, each coroutine in a graph processing task represents a unit of concurrent execution that processes the subgraph from the overall task. We can reduce the overhead of cache misses by overlapping the time of fetching data through the switching of different coroutines. However, switching coroutines also causes an overhead. Suppose a coroutine requires much more switch times than other coroutines. Under the scheduling by pooling scheduler (algorithm 1 in algorithm 1), it will occupy the most time of the algorithm’s execution with other coroutines having no overlapping with it. In such case, it will lead to an unbalanced interleaving execution mode, as illustrated in ”Vertex Table Partition” in Figure 7 3.
@@
To this end, two solutions can solve the issues, the first is to avoid the unbalanced partition by ensuring that each coroutine has approximately the same switch times, and another to optimize the pooling scheduler to avoid the repeated suspension and resumption of the remaining single coroutine after all other coroutines have been destroyed. Correspondingly, we propose two techniques to address the issue: a graph partition strategy to ensure partition evenness and a trimmed-polling scheduler with fine-grained checkpoints. ”Vertex Table Partition” is the basic graph partition strategy by dividing the vertex table into continuous chunks, which could maintain the natural locality of the graph data  (Zhu et al., 2016). This approach may also lead to unbalance in some extreme cases, as partitioning on the vertex table means that the smallest granularity of the partition is the edge block of a single vertex. For example, if a vertex is a super vertex (e.g., having $95\%$ of the edges), the approach may fail.
@@
A Fine-grained Graph Partition Strategy, on the other hand, controls the balance by partitioning GTChain into continuous sub-chains, as shown in Figure 7 2. Suppose there are $N$ coroutines and CBList contains $X$ edge structure blocks, then each coroutine will process a sub-chain (cut from GTChain) containing $X/N$ edge blocks. While this strategy achieves nearly perfect balance, it requires the participation of all the vertices and edges in the entire graph computation, just like the conditions for using GTChain. However, even though we evenly divide the data before computation, if certain edge blocks are not involved in the computation, it still cannot form a balanced interleaving execution mode at run time. Therefore, ”GTChain Partition” can only be used for scan_vertices() with scan_edges($v_{src}$). In summary, we employ two graph partition strategies in GastCoCo, and their specific usage will be detailed in subsection 6.1.
@@
A Trimmed-Polling Scheduler. Compared with the pooling scheduler, a more proper approach would be to pass the number of uncompleted coroutines to the resumed coroutine such that it can decide when to switch. If the coroutine is the only one that remains working, it would not issue a suspension operator, as shown in algorithm 3. On the contrary, this may introduce extra overhead because of the additional checks in practice.

### v4

@@ 3. Overview of GastCoCo
This paper proposes GastCoCo, an in-memory system for dynamic graph processing that employs an interleaved execution mode combining coroutines and software prefetching. GastCoCo focuses on graph storage and coroutine-based prefetch co-design, proposing a novel dynamic graph data structure friendly to prefetching techniques. Besides, GastCoCo comprises various task allocating, prefetching, and scheduling strategies to mitigate the overhead of cache misses, thereby delivering robust performance for different graph processing tasks across various hardware environments.
As shown in Figure 3, GastCoCo is primarily composed of three layers: storage layer, execution layer, and adaptation layer. Given a graph processing task (i.e., graph computation or graph update), user execution begins by calling GastCoCo API, as shown in Table 1.
GastCoCo API activates the dynamic graph processing engine, at which point the system configuration probe in the adaptation layer detects the hardware environment to configure the system parameters of GastCoCo, e.g., the numbers of coroutines per thread and hybrid prefetching strategy. To process tasks using software prefetching, graph data should be partitioned into several parts and allocated to a coroutine within a coroutine pool. All coroutines collaborate according to the scheduling strategy to produce the final results. The execution strategy tuner tailors the partition and scheduling strategy to suit the specific task to achieve better performance. The task allocator partitions the graph data according to the partition strategy and constructs the coroutine pool, while the coroutine scheduler schedules coroutines according to the scheduling strategy. Finally, the executor performs all read and write operations on the storage layer.
@@
LoadGraph (file_path) Load graph data files.
UpdateVertex (obj_v, update_mode) Add, delete, and modify vertices.
UpdateEdge (obj_e, update_mode) Add, delete, and modify edges.
BatchUpdate (update_content) Handle batch update tasks.
ProcessVertex (f, active) Process vertices task.
ProcessEdge (dense_f, sparse_f, active) Process edges task.
Execution Layer (§5). The execution layer includes the task allocator which breaks down a task into multiple subtasks and allocates them to coroutines, the coroutine scheduler which controls the interleaving execution and prefetching, and the executor which executes the computation tasks according to the scheduling strategy.
Adaptation Layer (§6). The adaptation layer includes two components: the system configuration probe and the execution strategy tuner. They utilize prefabricated programs to probe the optimal coroutine parameters (e.g., number of coroutines per thread) and the most suitable execution and prefetching strategy for the current tasks and runtime environment.
@@ 4.1. Overview
For dynamic graph processing tasks, we have the following expectations for an ideal data structure. It should: (1) maintain cache locality as possible to produce efficient graph computation; (2) be easy to dynamically adjust for frequent graph updates. To enhance the performance from both perspectives, a basic idea is to tailor as closely as possible to fit the hardware prefetching mechanisms and prepare for the implementation of software prefetching.
Figure 4 presents the proposed data structure CBList and we elaborate on its details as follows. For storing vertices, we employ an ID-map table and a vertex table (array of vertex structures), as shown in Figure 4 1 and 2, to map the vertex’s original ID to the logical ID (unique number identifier used in graph computation). A record contains multiple fields as presented in Figure 4 3, which includes two pointers to the neighbors (i.e., traversal pointer and update/query pointer), size (number of adjacent edges), delete flag, and the level. Here, we use a level variable to distinguish the structure of the vertex’s neighborhood. A value 00 represents using small chunks for edge storage, while other numbers indicate using B+ trees and the value represents the number of leaf nodes in the tree.
@@ 4.2. Update-Read Balanced Edge Storage
We enumerate the advantages of our Edge Storage design as follows. Firstly, we use independent edge structures (refer to ADJ-like structures in Figure 5) for each vertex to store its neighborhood to promote parallel updates of edges and facilitate different implementations like chunks and B+ trees for different nodes. In contrast, in CSR-like structures (e.g., PCSR (Wheatman and Xu, 2018) and Teseo (Leo and Boncz, 2021)), nodes are stored in identical and shared edge structure (refer to the CSR-like structures in Figure 5), and to update neighbors of one vertex may affect the vertex records and memory locations of other edges. Secondly, scan_edges($v_{src}$) accessed from Traversal pointer in a B+ tree can be considered as sequential data access on a linked list, which can effectively utilize hardware prefetching as discussed in subsection 2.2. Thirdly, read_edge($v_{src}$,$v_{dst}$) accessed from update/query pointer completes in $O(log(n))$.
Figure 5. (a) CSR-like physically contiguous store; (b) prefetch-friendly logically contiguous store via pointers; (c) ADJ-like mutually independent store (E.S.: Edge Storage).
@@ 4.3. Prefecth-Friendly Global Traversal Chain
In graph computation, a significant portion of tasks (e.g., PageRank (Page et al., 1998)) require all vertices and edges in the graph to be involved in the computation. Their data access patterns can be represented as a combination of scan_vertices() and scan_edges($v_{src}$). This combination equates to sequential data access over the entire graph.
@@
We have also considered the overhead of forming GTChain. One of the reasons we chose B+ trees is the ease of forming GTChain. Each B+ tree inherently contains a traversal chain formed by leaf nodes. Its implementation for GTChain simply involves using pointers to link the leaf node chain to the global chain. However, most tree structures (e.g., balanced binary trees, B-trees, and red-black trees) require manual implementation of a traversal chain and the additional traversal chain is challenging to form as a singly linked list (each node has only one outgoing pointer) to support the jump-pointer mechanism of hardware prefetching.
@@ 5.1. Coroutine with Software Prefetching
Sequential execution and interleaved execution. This paper focuses on reducing the cache miss overhead based on the interleaving execution mode (Psaropoulos et al., 2017). Figure 6 illustrates the difference between sequential execution and interleaving execution of traversal on a graph in Figure 6 1. We first demonstrate an example of scan_edges($v_{src}$) on AL (adjacency list) in Figure 6 2 for clarity: In the sequential execution mode, when the CPU accesses Node 1, the first neighbor of vertex A, a cache miss is likely to occur in the non-contiguous memory linked by pointers, making the CPU stalled, i.e., waiting until the data is fetched. Such stalling (He et al., 2020) occurs whenever cache misses happen due to non-contiguous memory access and pointer chasing. Alternatively, in interleaved execution mode, the CPU will issue a prefetch for Node 1 and switch to another task (i.e., accessing Node 3, the first neighbor of vertex B). Each coroutine handles a linked list and is in charge of computing the corresponding task. The algorithm starts from an arbitrary computation task. It switches to another task whenever it encounters a software prefetching instruction to fetch data (the instruction is set before accessing the next node in the linked list). When switching back to the previous task, the CPU does not need to wait for the needed data as they have been loaded into the cache. Thus, the process of fetching data and computations can be overlapped and the time of cache misses due to pointer chasing can be avoided. Note that switching tasks also causes time costs (Psaropoulos et al., 2017). We choose C++20 stackless coroutines and design scheduling strategies, such that the switching time is minimal, thereby achieving performance enhancements.
@@
Coroutine in Graph Computation. Next, we present the coroutine pool constructor and scheduler in graph computation. As we previously mentioned, C++20 stackless coroutines are a special kind of function and it requires manual intervention to schedule these functions. ①In algorithm 1 (algorithm 1 - algorithm 1), we present the basic logic of the coroutine pool constructor. The time complexity of the constructor is $O(m)$, where $m$ is the number of coroutines in the coroutine pool. The coroutine pool relies on the for loop (algorithm 1) to initialize each coroutine as Func. Func defines the graph access operations (as proposed in subsection 2.2) which will be detailed in algorithm 2. ②In algorithm 1 (algorithm 1 - algorithm 1), we present a polling scheduler implementation. The overhead generated by the scheduler itself is also a cost of using coroutines. Therefore, we design it as simply as possible to reduce the cost of using coroutines. Its core logic includes checking if the coroutine pool has finished through a while loop (algorithm 1) and polling each coroutine via a for loop (algorithm 1). If a coroutine is not completed, the polling scheduler resumes execution from the suspension point (algorithm 1). If a coroutine is completed, the polling scheduler will destroy it (algorithm 1). Assuming the number of suspensions for the $i$-th coroutine in the polling scheduler is $k_{i}$, the time complexity for the entire polling scheduler is $O(\sum_{i=1}^{m}k_{i})$. If the tasks are evenly distributed and each coroutine is suspended approximately $k$ times, the overall time complexity is $O(mk)$. ③We detail the logic of edge access operations in algorithm 2. GetNeighbors($vertex$) and GetNeighbors($chain$) correspond to the scan_edges($v_{src}$) operation for accessing edges and their usage and difference will be elaborated in subsection 6.1. The time complexities of the two operations are $O(D)$ and $O(E_{chain})$ respectively, where $D$ is the degree of the vertex, and $E_{chain}$ is the number of edges in the chain. FindNeighbor($edge$) corresponds to read_edge($v_{src}$,$v_{dst}$). Based on CBList, regardless of whether using a binary search on small chunks or searching within a B+ tree, the time complexity for FindNeighbor($edge$) is $O(\log D)$. The user can specify the search algorithm either on small chunks or B+ tree leaf nodes (refer to algorithm 2, algorithm 2). Here, all edge access operations are constructed using coroutines and prefetch($ptr$) denotes prefetching the content pointed to by the pointer in algorithm 2, algorithm 2, and algorithm 2. This is in the form of software prefetching. In algorithm 2, algorithm 2, and algorithm 2, co_await is a keyword in C++20 and its invocation of suspend_always means that this coroutine actively entering an always suspended state. Coroutines in the suspended state will be resumed by the scheduler (algorithm 1 in algorithm 1).
@@ 8 Function Scheduler($coroutine\_pool$):
15                         $coroutine\_pool[i].$destroy()
@@ 5.2. Load Balancing of Coroutines
As we have described in the example of interleaving execution mode in Figure 6, each coroutine in a graph processing task represents a unit of concurrent execution that processes the subgraph from the overall task. We can reduce the overhead of cache misses by overlapping the time of fetching data through the switching of different coroutines. However, switching coroutines also causes an overhead. Suppose a coroutine requires much more switch times than other coroutines. Under the scheduling by polling scheduler (algorithm 1 in algorithm 1), it will occupy the most time of the algorithm’s execution, with other coroutines having no overlapping with it. In such case, it will lead to an unbalanced interleaving execution mode, as illustrated in “Vertex Table Partition” in Figure 7 3.
@@
To this end, two solutions can solve the issues. The first is to avoid the unbalanced partition by ensuring that each coroutine has approximately the same switch times, and another is to optimize the polling scheduler to avoid the repeated suspension and resumption of the remaining single coroutine after all other coroutines have been destroyed. Correspondingly, we propose two techniques to address the issue: a graph partition strategy to ensure partition evenness and a trimmed-polling scheduler with fine-grained checkpoints. “Vertex Table Partition” is the basic graph partition strategy by dividing the vertex table into continuous chunks, which could maintain the natural locality of the graph data  (Zhu et al., 2016). This approach may also lead to unbalance in some extreme cases, as partitioning on the vertex table means that the smallest granularity of the partition is the edge block of a single vertex. For example, if a vertex is a super vertex (e.g., having $95\%$ of the edges), the approach may fail.
@@
A Fine-grained Graph Partition Strategy, on the other hand, controls the balance by partitioning GTChain into continuous sub-chains, as shown in Figure 7 2. Suppose there are $N$ coroutines and CBList contains $X$ edge structure blocks, then each coroutine will process a sub-chain (cut from GTChain) containing $X/N$ edge blocks. While this strategy achieves nearly perfect balance, it requires the participation of all the vertices and edges in the entire graph computation, just like the conditions for using GTChain. However, even though we evenly divide the data before computation, if certain edge blocks are not involved in the computation, it still cannot form a balanced interleaving execution mode at run time. As a result, “GTChain Partition” can only be used for scan_vertices() with scan_edges($v_{src}$). In summary, we design and employ two graph partition strategies in GastCoCo, and their specific usage will be detailed in subsection 6.1.
@@
A Trimmed-Polling Scheduler. Compared with the polling scheduler, a more proper approach would be to pass the number of uncompleted coroutines to the resumed coroutine such that it can decide when to switch. If the coroutine is the only one that remains working, it would not issue a suspension operator, as shown in algorithm 3. On the contrary, this may introduce extra overhead because of the additional checks in practice.